The height in feet of a glof ball hit from an elevated tee box is given by h(t) = -16t^2 + 60t + 25 where t represents the time in seconds the ball is in flight.
determine the time interval for which the ball is above 25 ft.

Answer by Alan3354(69443)

The height in feet of a glof ball hit from an elevated tee box is given by h(t) = -16t^2 + 60t + 25 where t represents the time in seconds the ball is in flight.
determine the time interval for which the ball is above 25 ft.
------------
It's at 25 feet at t = 0
h(t) = -16t^2 + 60t + 25
25 = -16t^2 + 60t + 25
16t^2 = 60t
t = 60/16
t = 15/4 seconds to return to 25 ft
The time above 25 ft is
15/4>t>0 seconds